Circular No. 8326

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 F4 (BRADFIELD) Z. Sekanina, Jet Propulsion Laboratory, writes: “The comet’s peculiar

shape in the SOHO’s C3 coronagraphic images is caused by two effects. The gradual coma broadening in a direction nearly perpendicular to the apparent motion, beginning before Apr. 18.8 UT, is caused by a rapid development of the comet’s post-perihelion dust tail, which points to the southwest from the nucleus in the images taken around Apr. 19.5 and to the west in the more recent C3 images. The increasingly rugged boundary of the tail on its antisolar side, especially well pronounced after Apr. 18.8, is due to discrete features protruding through the tail’s smooth leading edge. These protrusions are either streamers — i.e., synchronic formations that are products of recurring events of briefly elevated dust emission (outbursts) — or striae (i.e., pseudosynchronic formations of a more complex nature). If they are streamers (perhaps a more likely case, judging from their appearance in the images just before the comet’s head left the C3 field-of-view), their sequence suggests fairly consistently an outburst recurrence period of ∼ 0.5 day, which might indicate the comet’s rotation period. For example, in the image of Apr. 19.496, one can identify at least seven features, whose position angles are estimated at between 150^{o} and 240^{o}, with the corresponding calculated event times between 2.4 days before perihelion and 0.6 day after perihelion. The peak radiation-pressure accelerations on the particles in the features are estimated at ∼ 2–3 times the sun’s gravitational acceleration.”

K. Battams, Naval Research Laboratory, has reported measurements of the position of C/2004 F4 from the SOHO C3 images. The reduced astrometry by B. G. Marsden and R. Kisala and the following orbital elements appear on MPEC 2004-H37.

T = 2004 Apr. 17.129 TT  ω = 333.^{o}298
Ω = 222.992 }2000.0
q = 0.16749 AU  i = 63.118
2004TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Apr. 20 1^{h}24.^{m}29 +14^{o}20.^{′}3 0.837 0.205 7.^{o}4 140.^{o}9  2.5
25 1 03.97 +25 36.5 0.938 0.345 20.1 91.1  4.4
30 0 59.13 +32 24.5 1.067 0.491 27.2 69.6  5.8
May 5 1 00.16 +37 00.6 1.190 0.627 31.8 58.0  6.9
10 1 03.54 +40 28.2 1.302 0.754 35.3 50.7  7.7
15 1 07.85 +43 15.8 1.405 0.874 38.2 45.7  8.3
2004 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green